Solve each system by the addition method.
2x + 4y = 28
x - 2y = 10

Answer by nerdybill(7384)

2x + 4y = 28
x - 2y = 10
.
Multiply bottom equation by 2 and combine both equations:
2x + 4y = 28
2x - 4y = 20
---------------
4x = 48
x = 12
.
plug above back into:
x - 2y = 10
12 - 2y = 10
-2y = -2
y = 1
.
Solution: (x,y) = (12, 1)